OpenStudy (anonymous):

express as a single log 2/3logbase2 8 + 1/3 logbase2 3 - 2logbase2 4

OpenStudy (dumbcow):

use log rules n*logx = log(x^n) logx +logy = log(x*y) logx -logy = log(x/y)

OpenStudy (anonymous):

i know te rules, want to verify my work done

OpenStudy (dumbcow):

oh ok \[=\log_{2}\frac{8^{2/3}*3^{1/3}}{16}\rightarrow \log_{2}\frac{3^{1/3}}{4} \]

OpenStudy (anonymous):

thanks